# No Need to Look Back: An Efficient and Scalable Approach for Temporal Network Representation Learning

### v1

@@ 4.4 Node representation generation and NLB prediction
$\text{where}\,\{\alpha_{i}\}=\text{\small{softmax}}(\{w^{T}\text{MLP}(r_{v}^{t},e^{t^{\prime}}_{u,v},t^{\prime})|(v,e^{t^{\prime}}_{u,v},t^{\prime})\in S^{t}_{u}\})$, where $w$ is a learnable vector parameter. A pair of node representations $Z^{t}_{u}$ and $Z^{t}_{v}$ can be combined and plugged in to classifiers to make link predictions; and a single node representation $Z^{t}_{u}$ can be used for node classification. $t$ is encoded via T-encoding (see Def. C.1 in Appendix C) which has been proved to be effective in previous works (Xu et al., 2019; Kazemi et al., 2019a; Xu et al., 2020; Wang et al., 2021d).
Note that the node status $r_{u}^{t}$ will get updated via $r_{u}^{t+1}\leftarrow\textbf{RNN}([r_{u}^{t},r_{v}^{t},t,e^{t}_{u,v}])$ after an event $(u,v,t)$ arrives.
Also, NLB naturally supports efficient aggregation of second-hop temporal neighbors because to get representation for node $u$, if $v$ is stored in $S_{u}^{t}$, i.e., one of the first-hop temporal neighbors, $S_{v}^{t}$ provides the down-sampled second-hop neighbors of $u$, which can be efficiently accessed. However, in experiment, we use the first hop only as it already provides decent prediction performance. Involving the second hop, albeit bringing more information, will somewhat increase the inference latency.
@@ 5 Experiments
Measurement Wikipedia Reddit GDELT MAG Ubuntu Wiki-talk
nodes 9K 11K 17K 122M 159K 1.1M
temporal links 157K 672K 191M 1.3B 964K 7.8M
node classes 2 2 81 152 0 0
labels 217 366 42M 1.4M 0 0
node features 0 0 413 768 0 0
edge features 172 172 186 0 0 0
Table 2: Summary of dataset statistics.
@@
5.1 Experimental setup
Datasets. We use six publicly available real-world datasets whose statistics are listed in Table 2 for experiments. There are datasets with billion-scale temporal links and million-scale nodes. Further details of these datasets can be found in Appendix D. We split the datasets into training, validation and testing data according to the ratio of 70/15/15.
Baselines. We select five representative baselines. Among the baselines, DySAT (Sankar et al., 2020) is a snapshot-based approach, TGAT (Xu et al., 2020) proposed aggregating temporal neighborhood information with attention, TGN (Rossi et al., 2020) proposed keeping a memory state for each node that gets update with new interactions, APAN (Wang et al., 2021c) proposed asynchronous graph propagation instead of graph aggregation, and NAT (Luo & Li, 2022) improves link prediction with their joint neighborhood structural features. For DySAT, TGAT, TGN and APAN, we evaluate their performance and scalability with two sampling strategies: uniform sampling (denoted as unif) and truncation of the most recent neighbors (denoted as trunc). We evaluate these baselines (excluding NAT) using the implementation in TGL (Zhou et al., 2022) that relies on low-level C++ programming and multi-thread CPU, as it is currently the most efficient framework for them and the original implementation has shown to be significantly slower. Additional details about the baselines can be found in Appendix E.
Hyperparameters. For fair comparison, we fix the maximum number of neighbors to be sampled and the training and inference batch sizes for each dataset across all methods which is specified in Table 8 (in Appendix). We limit the maximum number of CPU threads to be at most 32 and we use one GPU for all methods. For the rest of the hyperparameters, if a dataset has been tested previously, we use the set of hyperparameters that are provided by the baseline models. Otherwise, we tune the parameters and make sure the sizes of different modules such as the node representation and time feature, have dimensions of the same scale for all methods. More detailed hyperparameters are provided in Appendix E. Hardwares. We run all experiments using the same Linux device that is equipped with 256 AMD EPYC 7763 64-Core Processor @ 2.44GHz with 2038GiB RAM and one GPU (NVIDIA RTX A6000).
@@ 5.1 Experimental setup
Method Train Test Inf. Lat. CPU (MJ) GPU (MJ)
Ubuntu TGN-trunc 15.95 2.05 1.37 3,332 1.75
TGN-unif 15.55 2.11 1.42 3,326 1.78
TGAT-trunc 12.37 1.70 1.20 3,028 1.30
TGAT-unif 13.64 1.93 1.27 3,254 1.37
APAN-trunc 12.96 1.94 0.61 2,802 1.36
APAN-unif 13.01 2.07 0.62 2,894 1.34
DySAT-trunc 23.51 2.79 2.34 4,442 1.79
DySAT-unif 23.89 2.88 2.40 4,327 1.75
NAT 17.52 1.96 1.22 2,853 1.27
NLB-edge 11.06 1.30 0.57 2,764 2.24
NLB-node 10.74 1.35 0.59 2,748 2.17
Method Train Test Inf. Lat. CPU (MJ) GPU (MJ)
Wiki-talk TGN-trunc 74.63 11.76 7.61 16,715 9.57
TGN-unif 85.96 13.77 8.19 18,925 11.2
TGAT-trunc 58.37 8.74 5.88 14,787 5.90
TGAT-unif 63.32 9.14 6.35 16,553 6.75
APAN-trunc 60.68 10.34 2.98 13,898 6.94
APAN-unif 64.06 10.61 2.88 14,463 7.04
DySAT-trunc 113.71 14.94 12.16 22,572 9.28
DySAT-unif 115.27 15.29 12.56 22,889 9.24
NAT 81.32 10.03 6.08 14,736 7.13
NLB-edge 59.18 7.30 3.03 12,277 15.0
NLB-node 56.36 7.01 2.99 11,519 13.6
Method Train Test Inf. Lat. CPU (MJ) GPU (MJ)
GDELT TGN-trunc 1977 406 302 431,567 174.6
TGN-unif 4791 1013 777 797,311 389.4
TGAT-trunc 1123 248 183 306,860 92.7
TGAT-unif 1493 383 301 251,452 83.8
APAN-trunc 1530 338 200 485,818 211.2
APAN-unif 1839 347 187 512,061 188.4
DySAT-trunc 2696 535 437 350,263 107.3
DySAT-unif 4764 907 749 526,472 185.1
NAT - - - - -
NLB-edge 1323 139 63 252,267 307.6
NLB-node 1089 118 60 215,820 258.5
Method Train Test Inf. Lat. CPU (GJ) GPU (GJ)
MAG TGN-trunc 6845 1266 795 15,041 1.59
TGN-unif 5704 1224 741 21,334 2.02
TGAT-trunc 3465 754 518 10,427 1.06
TGAT-unif 3520 774 541 13,236 1.35
APAN-trunc 12347 1600 809 22,262 2.45
APAN-unif - - - - -
DySAT-trunc 6195 1469 1213 14,975 1.48
DySAT-unif 5587 1562 1291 10,786 1.25
NAT - - - - -
NLB-edge 3073 593 296 2,684 1.73
NLB-node 2729 577 279 2,870 1.69
Table 5: Scalability evaluation on all datasets. Note that MJ = $10^{6}$ Joules, and GJ = $10^{9}$ Joules.
Performance Metrics. For link prediction performance, we evaluation all models with Area Under the ROC curve (AUC) and Average Precision (AP). In the main text, the prediction performance in all tables is evaluated in AUC. For node classification, we evaluate with AUC for Wikipedia and Reddit which have two node classes, and we use F1 for GDELT and MAG which have more than two classes. All results are summarized based on 5 time independent experiments.
@@ 5.2 Results and Discussion
Prediction performance. The link prediction performance in AUC is given in Table 3 and the node classification performance is given in Table 4. The link prediction result in AP is given in Appendix Table 7.
For link prediction, NLB-edge achieves either the best or the second best performances in 11 out of 12 settings (6 datasets with inductive and transductive tasks), while NLB-node closely follows or sometimes marginally surpasses NLB-edge. NAT is the best performing baseline for link prediction because it leverages the joint neighborhood structural features (Li et al., 2020; Wang et al., 2021d) designed for node-pairs. NAT cannot be applied to node classification. Moreover, NAT consumes significant GPU memory. TGN-trunc and TGN-unif are the second most competitive baselines. They achieve similar performance to NLB-edge and NLB-node sometimes but significantly fall behind on the datasets Ubuntu and Wiki-talk, and 3 of the 4 node classification tasks. One aspect that differentiate Ubuntu and Wiki-talk to the rest of the dataset is that they do not have node or edge features (Table 2). We hypothesize that the benefit of recent sampling is more obvious with the absence of additional features because NLB can gather effective historical information from diverse and recent temporal neighbors. The other baselines usually perform far behind NLB.
For node classification, both NLB-edge and NLB-node achieve similar or better results than the baselines. In MAG, all baselines fail to perform well and we cannot reproduce the scores reported in Zhou et al. (2022) (more discussion in Appendix F).
NLB-edge often outperforms NLB-node in prediction performance because NLB-edge tends to encode more temporal neighbors (node IDs plus timestamps) while NLB-node only encodes neighbors with unique node IDs.
@@
On inference latency, except for APAN on the Wiki-talk dataset, both NLB-node and NLB-edge have shown significant improvements over all settings. APAN aims to improve inference latency but consumes massive CPU memory. On large-scale datasets such as GDELT and MAG, APAN requires significant latency for transfering messages from CPU to GPU and no longer has advantages. For all other baselines than APAN, NLB-node gives 1.97-5.02$\times$ speedup on inference latency over all datasets.
Although NLB’s GPU energy consumption is slightly larger, it is negligible compared to the CPU energy consumption. Compared to NLB-edge, NAT consumes up to 1.8$\times$ more energy while TGN-trunc and TGN-unif consumes 1.2-5.6$\times$ and 1.2-7.94$\times$ more energy over all datasets respectively. On billion-scale datasets MAG, NLB-edge is 3.88-8.29$\times$ more energy efficient than all methods.
@@
We also plot the model convergence curves in Fig. 2 on the link prediction validation performance v.s. CPU/GPU wall-clock training time on Reddit. NLB-edge and NLB-node are among the fastest to converge to a high performance.
@@ 5.3 Additional analysis
Figure 3: The changes in transductive link prediction test performance on Reddit and in inductive link prediction test performance on Ubuntu with respect to different $\alpha$.
s Transductive Inductive Node Class. Train Test Inf. Lat. CPU (MJ) GPU (MJ)
@@
Table 6: The performance and scalability of NLB-edge and NLB-node on the Reddit dataset given different table sizes $s$, the maximum number of temporal neighbors sampled. $s=20$ is the setting used for comparison with baselines.
6 Conclusion and Future Works
In this work, we introduced NLB which abandons the traditional time-consuming temporal neighbor backtracking and sampling while adopting the newly proposed forward recent sampling. We proved that NLB-edge achieves recent sampling of links and NLB-node achieves recent sampling of nodes. Our extensive experiments demonstrate that NLB is effective in both prediction performance and scalability while significantly improving inference latency and energy efficiency. In the future, we plan to extend NLB to even larger-scale networks whose historical information of nodes and edges has to be stored in distributed systems.
@@ 7 Acknowledgement
The authors would like to thank Hongkuan Zhou of USC for the time to discuss the code for TGL (Zhou et al., 2022). This project is partially supported by NSF awards OAC-2117997, IIS-2239565, and 2023 JPMC faculty award.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix C T-encoding
It has shown to be effective for TGRL (Xu et al., 2019; Kazemi et al., 2019a; Xu et al., 2020; Wang et al., 2021d; Rossi et al., 2020; Liu et al., 2022).
@@
Transductive TGN-truncate 99.44 $\pm$ 0.08 99.59 $\pm$ 0.09 98.34 $\pm$ 0.10 99.30 $\pm$ 0.06 81.61 $\pm$ 0.20 87.03 $\pm$ 2.47
TGN-uniform 99.36 $\pm$ 0.04 99.60 $\pm$ 0.05 96.96 $\pm$ 0.05 99.21 $\pm$ 0.04 83.27 $\pm$ 0.96 87.24 $\pm$ 0.16
TGAT-truncate 97.65 $\pm$ 0.04 97.98 $\pm$ 0.03 98.12 $\pm$ 0.01 99.08 $\pm$ 0.03 80.62 $\pm$ 0.39 84.81 $\pm$ 0.06
TGAT-uniform 94.61 $\pm$ 0.23 98.12 $\pm$ 0.15 97.54 $\pm$ 0.01 99.01 $\pm$ 0.05 81.82 $\pm$ 0.12 84.94 $\pm$ 0.01
APAN-truncate 99.04 $\pm$ 0.03 95.73 $\pm$ 2.96 96.79 $\pm$ 0.32 87.85 $\pm$ 1.68 72.20 $\pm$ 3.94 84.13 $\pm$ 4.44
APAN-uniform 97.14 $\pm$ 0.67 96.57 $\pm$ 0.24 96.21 $\pm$ 1.35 CPU OOM 78.61 $\pm$ 1.61 85.68 $\pm$ 6.00
DySAT-truncate 96.10 $\pm$ 0.02 97.96 $\pm$ 0.01 96.03 $\pm$ 0.27 98.97 $\pm$ 0.10 63.93 $\pm$ 0.01 72.47 $\pm$ 0.02
DySAT-uniform 96.40 $\pm$ 0.04 97.98 $\pm$ 0.02 98.45 $\pm$ 0.08 98.95 $\pm$ 0.35 64.24 $\pm$ 0.01 74.64 $\pm$ 0.01
@@
Inductive TGN-truncate 98.42 $\pm$ 0.08 99.38 $\pm$ 0.07 98.38 $\pm$ 0.04 96.99 $\pm$ 0.08 82.27 $\pm$ 1.17 88.46 $\pm$ 0.64
TGN-uniform 97.99 $\pm$ 0.13 99.21 $\pm$ 0.03 98.48 $\pm$ 0.15 97.21 $\pm$ 0.14 85.70 $\pm$ 1.58 88.10 $\pm$ 0.13
TGAT-truncate 96.75 $\pm$ 0.02 95.77 $\pm$ 0.21 94.64 $\pm$ 0.01 98.78 $\pm$ 0.01 81.53 $\pm$ 0.14 84.10 $\pm$ 0.05
TGAT-uniform 92.94 $\pm$ 0.15 96.33 $\pm$ 0.09 93.32 $\pm$ 0.02 98.75 $\pm$ 0.01 82.17 $\pm$ 0.08 82.21 $\pm$ 0.05
APAN-truncate 95.18 $\pm$ 0.91 96.02 $\pm$ 2.06 97.99 $\pm$ 0.03 CPU OOM 68.60 $\pm$ 2.41 80.47 $\pm$ 0.35
APAN-uniform 95.73 $\pm$ 0.77 96.03 $\pm$ 1.31 96.55 $\pm$ 0.31 CPU OOM 78.64 $\pm$ 2.66 88.82 $\pm$ 0.28
DySAT-truncate 95.25 $\pm$ 0.22 97.34 $\pm$ 0.04 89.58 $\pm$ 0.06 98.76 $\pm$ 0.01 64.50 $\pm$ 0.15 72.06 $\pm$ 0.33
DySAT-uniform 95.26 $\pm$ 0.06 97.47 $\pm$ 0.07 94.14 $\pm$ 0.22 98.74 $\pm$ 0.01 64.78 $\pm$ 0.05 74.12 $\pm$ 0.26
@@
Appendix D Dataset Description
@@
  • Wikipedia [^2] http://snap.stanford.edu/jodie/wikipedia.csv records edit events on wiki pages. It is a bipartite graph where a set of nodes represents the editors and another set represents the wiki pages. The stamped link represents the edit events. The edge features are extracted from the contents of wiki pages. The node labels are binary which indicates whether a user is banned from posting.
@@
  • Reddit333 http://snap.stanford.edu/jodie/reddit.csv is a dataset of the post events by users on subreddits. It is also an attributed bipartite graph between users and subreddits. The node labels are also binary values that indicate whether a user is banned from posting to a subreddit.

### v2

@@ 4.4 Node representation generation and NLB prediction
$\text{where}\,\{\alpha_{i}\}=\text{{softmax}}(\{w^{T}\text{MLP}(r_{v}^{t},e^{t^{\prime}}_{u,v},t^{\prime})|(v,e^{t^{\prime}}_{u,v},t^{\prime})\in S^{t}_{u}\})$, where $w$ is a learnable vector parameter. A pair of node representations $Z^{t}_{u}$ and $Z^{t}_{v}$ can be combined and plugged in to classifiers to make link predictions; and a single node representation $Z^{t}_{u}$ can be used for node classification. $t$ is encoded via T-encoding (see Def. C.1 in Appendix C) which has been proved to be effective in previous works [46, 47, 17, 16].
Note that a node status $r_{u}^{t}$ is updated via $r_{u}^{t+1}\leftarrow\textbf{RNN}([r_{u}^{t},r_{v}^{t},t,e^{t}_{u,v}])$ after an event $(u,v,t)$ arrives.
Also, NLB naturally supports efficient aggregation of second-hop temporal neighbors because to get representation for node $u$, if $v$ is stored in $S_{u}^{t}$, i.e., one of the first-hop temporal neighbors, $S_{v}^{t}$ provides the down-sampled second-hop neighbors of $u$, which can be efficiently accessed. However, in experiment, we use the first hop only as it already provides decent prediction performance. Involving the second hop, albeit bringing more information, will increase the inference latency.
@@ 5 Experiments
5.1 Experimental setup
Datasets. We use six publicly available real-world datasets whose statistics are listed in Appendix Table 6 for experiments. There are datasets with billion-scale temporal links and million-scale nodes. Further details of these datasets can be found in Appendix D. We split the datasets into training, validation and testing data according to the ratio of 70/15/15 while preserving the chronological order.
Baselines. We select four representative baselines, each with two variants. Among the baselines, TGAT [17] proposed aggregating temporal neighborhood information with attention, TGN [15] proposed keeping a memory state for each node that gets update with new interactions, APAN [21] proposed asynchronous graph propagation instead of graph aggregation, and NAT [18] improves link prediction with their joint neighborhood structural features. For TGAT, TGN and APAN, we evaluate their performance and scalability with two sampling strategies: uniform sampling (denoted as unif) and truncation of the most recent neighbors (denoted as trunc). We evaluate these baselines (excluding NAT) using the implementation in TGL [25] that relies on low-level C++ programming and multi-thread CPU, as it is currently the most efficient framework for them and the original implementation has shown to be significantly slower. For NAT, we adapt it for node classification (named NAT-node) by removing their joint structural features and directly aggregating their neighborhood representations to generate node representations. Additional details about these baselines can be found in Appendix E.
Hyperparameters. For fair comparison, we fix the maximum number of neighbors to be sampled for each dataset across all methods as specified in Appendix Table 9. When evaluating scalability, we (1) limit the maximum number of CPU threads to be at most 32, (2) use one GPU for all methods and (3) fix the batch sizes across different methods as specified in Appendix Table 10. For the rest of the hyperparameters, if a dataset has been tested previously, we use the set of hyperparameters that are provided by the baseline models. Otherwise, we tune the parameters with grid search and make sure the sizes of different modules such as the node representation and time feature, have dimensions of the same scale for all methods. More detailed hyperparameters are provided in Appendix E. Hardwares. We run all experiments using the same Linux device that is equipped with 256 AMD EPYC 7763 64-Core Processor @ 2.44GHz with 2038GiB RAM and one GPU (NVIDIA RTX A6000).
Performance Metrics. For link prediction performance, we evaluation all models with Area Under the ROC curve (AUC), Average Precision (AP) and Mean Reciprocal Rank (MRR) with a large number of negative samples per positive sample (50 for GDELT, and 500 for other datasets, excluding the MAG dataset due to the time and memory constraints). In the main text, the prediction performance in all tables is evaluated in AUC. For node classification, we evaluate with AUC for Wikipedia and Reddit which have two node classes, and we use F1 for GDELT and MAG which have more than two classes. All results are summarized based on 5 time independent experiments with different random seeds and initializations.
@@ 5.1 Experimental setup
Train Test Inf. Lat. CPU (MJ) GPU (MJ)
Ubuntu 15.95 2.05 1.37 3,332 1.75
15.55 2.11 1.42 3,326 1.78
12.37 1.70 1.20 3,028 1.30
13.64 1.93 1.27 3,254 1.37
12.96 1.94 0.61 2,802 1.36
13.01 2.07 0.62 2,894 1.34
17.52 1.96 1.22 2,853 1.27
15.60 1.60 1.30 2,824 1.23
11.06 1.30 0.57 2,764 2.24
10.74 1.35 0.59 2,748 2.17
Train Test Inf. Lat. CPU (MJ) GPU (MJ)
Wiki-talk 74.63 11.76 7.61 16,715 9.57
85.96 13.77 8.19 18,925 11.2
58.37 8.74 5.88 14,787 5.90
63.32 9.14 6.35 16,553 6.75
60.68 10.34 2.98 13,898 6.94
64.06 10.61 2.88 14,463 7.04
81.32 10.03 6.08 14,736 7.13
67.28 8.79 4.88 14,492 6.82
59.18 7.30 3.03 12,277 15.0
56.36 7.01 2.99 11,519 13.6
Train Test Inf. Lat. CPU (MJ) GPU (MJ)
GDELT 1977 406 302 431,567 174.6
4791 1013 777 797,311 389.4
1123 248 183 306,860 92.7
1493 383 301 251,452 83.8
1530 338 200 485,818 211.2
1839 347 187 512,061 188.4
1323 139 63 252,267 307.6
1089 118 60 215,820 258.5
Train Test Inf. Lat. CPU (GJ) GPU (GJ)
MAG 6845 1266 795 15,041 1.59
5704 1224 741 21,334 2.02
3465 754 518 10,427 1.06
3520 774 541 13,236 1.35
12347 1600 809 22,262 2.45
3073 593 296 2,684 1.73
2729 577 279 2,870 1.69
Table 4: Scalability evaluation on all datasets. Note that MJ = $10^{6}$ Joules, and GJ = $10^{9}$ Joules. Note that TGL is adopted to implement baselines including TGN, TGAT, and APAN.
@@ 5.2 Results and Discussion
Prediction performance. The node classification performance is given in Table 3 and the link prediction performance in AUC is given in Table 2. The link prediction results in AP and in MRR are given in Appendix Table 7 and Table 8 respectively.
For link prediction, NLB-edge achieves either the best or the second best performances in 10 out of 12 settings (6 datasets with inductive and transductive tasks), while NLB-node closely follows or sometimes marginally surpasses NLB-edge. NAT is the best performing baseline for link prediction because it leverages the joint neighborhood structural features [48, 16] designed for node-pairs. However, NAT cannot be applied to node classification, and it consumes significant GPU memory. We discuss NAT-node in detail later in this section. TGN-trunc and TGN-unif are the second most competitive baselines. They achieve similar performance to NLB-edge and NLB-node sometimes but significantly fall behind on the datasets Ubuntu and Wiki-talk, and 3 of the 4 node classification tasks. One aspect that differentiate Ubuntu and Wiki-talk to the rest of the dataset is that they do not have node or edge features (Table 6). We hypothesize that the benefit of recent sampling is more obvious with the absence of additional features because NLB can gather effective historical information from diverse and recent temporal neighbors. The other baselines usually perform far behind NLB.
For node classification, both NLB-edge and NLB-node achieve similar or better results than the baselines. In MAG, all baselines fail to perform well and we cannot reproduce the scores reported in Zhou et al. [25] (more discussion in Appendix G).
NLB-edge often outperforms NLB-node because NLB-edge tends to encode more temporal neighbors (node IDs plus timestamps) while NLB-node only encodes neighbors with unique node IDs.
@@
On inference latency, except for APAN on the Wiki-talk dataset, both NLB-node and NLB-edge shows significant improvements over all settings. APAN aims to improve inference latency but consumes massive CPU memory. On large-scale datasets such as GDELT and MAG, APAN requires significant latency for transfering messages from CPU to GPU and no longer has advantages. Apart from APAN, NLB-node gives 1.63-12.95$\times$ speedup on inference latency over all datasets.
Although NLB’s GPU energy consumption is slightly larger, it is negligible compared to the CPU energy consumption. Compared to NLB-edge, NAT consumes up to 1.8$\times$ more energy while TGN-trunc and TGN-unif consumes 1.2-5.6$\times$ and 1.2-7.94$\times$ more energy over all datasets respectively. On billion-scale datasets MAG, NLB-edge is 3.88-8.29$\times$ more energy efficient than all methods. The large difference between the energy consumed by the CPU v.s. the GPU can be attributed to (1) the default energy costs by CPU when no application is running, and (2) the intensive CPU-GPU communications which is counted toward the CPU cost as it controls the communication (more details in Appendix H).
@@
We further compare the efficiency of different sampling approaches in Appendix I, which shows that forward sampling is significantly faster. We also plot the model convergence curves in Fig. 2 on the link prediction validation performance v.s. CPU/GPU wall-clock training time on Reddit. NLB-edge and NLB-node are among the fastest to converge to a high performance.
Comparisons between NLB and NAT-node. Overall, NAT-node shows improvement over NAT on all of the scalability metrics. However, it does not outperform NLB in either computational efficiency or prediction performance. As shown by Table 4, over all datasets, NAT-node is still slower than NLB-edge (1.14-1.48$\times$ slower in training and 1.61-1.91$\times$ slower in inference). It still causes GPU out-of-memory error for large-scale datasets (GDELT and MAG). We think the computational inefficiency may be caused by the storage and maintenance of their neighborhood representations which are primarily used for structural features construction. Furthermore, NAT-node does not perform as well as NLB-edge on link prediction (Table 2) and node classification (Table 3).
@@ 5.3 Additional analysis
We give the evaluation of MRR for various $\alpha$’s and $s$’ in Appendix Table 13 and Table 14 respectively.
Figure 3: The changes in transductive link prediction test performance on Reddit (Left) and in inductive link prediction test performance on Ubuntu (Right) with respect to $\alpha$’s.
s Transductive Inductive Node Class. Train Test Inf. Lat. CPU GPU
@@
Table 5: The performance and scalability of NLB-edge and NLB-node on the Reddit dataset given different down-sampled neighbor hash table sizes $s$. $s=20$ is the setting used for comparison with baselines. CPU and GPU energies are measured in MJ.
6 Limitations and Future Work
We are aware of the limitation of NLB and highlight some future research directions. For industry-level graphs, it may not be possible to maintain the down-sampled temporal neighbors of all nodes in a single GPU. In the future, we plan to adopt distributed GPUs by partitioning the down-sampled neighbors of different nodes into different GPUs. We will guarantee that the memory used for each node is contained within the same GPU so that the maintenance can still be fast.
7 Conclusion
In this work, we introduced NLB which abandons the traditional time-consuming temporal neighbor backtracking and sampling while adopting the newly proposed forward recent sampling. We proved that NLB-edge achieves recent sampling of links and NLB-node achieves recent sampling of nodes. Our extensive experiments demonstrate that NLB is effective in both prediction performance and scalability while significantly improving inference latency and energy efficiency.
@@ 8 Acknowledgement
The authors would like to thank Hongkuan Zhou of USC for the time to discuss the code for TGL [25]. We would also like to thank all reviewers for providing valuable feedback for improving this work. Y. Luo and P. Li are supported by NSF awards PHY-2117997, IIS-2239565.
@@ Appendix C T-encoding
It has shown to be effective for TGRL [46, 47, 17, 16, 15, 49].
Appendix D Dataset Description
Measurement Wikipedia Reddit GDELT MAG Ubuntu Wiki-talk
nodes 9K 11K 17K 122M 159K 1.1M
temporal links 157K 672K 191M 1.3B 964K 7.8M
node classes 2 2 81 152 0 0
labels 217 366 42M 1.4M 0 0
node features 0 0 413 768 0 0
edge features 172 172 186 0 0 0
Table 6: Summary of dataset statistics.
@@ Appendix D Dataset Description
Transductive TGN-trunc 99.44 $\pm$ 0.08 99.59 $\pm$ 0.09 98.34 $\pm$ 0.10 99.30 $\pm$ 0.06 81.61 $\pm$ 0.20 87.03 $\pm$ 2.47
TGN-unif 99.36 $\pm$ 0.04 99.60 $\pm$ 0.05 96.96 $\pm$ 0.05 99.21 $\pm$ 0.04 83.27 $\pm$ 0.96 87.24 $\pm$ 0.16
TGAT-trunc 97.65 $\pm$ 0.04 97.98 $\pm$ 0.03 98.12 $\pm$ 0.01 99.08 $\pm$ 0.03 80.62 $\pm$ 0.39 84.81 $\pm$ 0.06
TGAT-unif 94.61 $\pm$ 0.23 98.12 $\pm$ 0.15 97.54 $\pm$ 0.01 99.01 $\pm$ 0.05 81.82 $\pm$ 0.12 84.94 $\pm$ 0.01
APAN-trunc 99.04 $\pm$ 0.03 95.73 $\pm$ 2.96 96.79 $\pm$ 0.32 87.85 $\pm$ 1.68 72.20 $\pm$ 3.94 84.13 $\pm$ 4.44
APAN-unif 97.14 $\pm$ 0.67 96.57 $\pm$ 0.24 96.21 $\pm$ 1.35 CPU OOM 78.61 $\pm$ 1.61 85.68 $\pm$ 6.00
@@
Inductive TGN-trunc 98.42 $\pm$ 0.08 99.38 $\pm$ 0.07 98.38 $\pm$ 0.04 96.99 $\pm$ 0.08 82.27 $\pm$ 1.17 88.46 $\pm$ 0.64
TGN-unif 97.99 $\pm$ 0.13 99.21 $\pm$ 0.03 98.48 $\pm$ 0.15 97.21 $\pm$ 0.14 85.70 $\pm$ 1.58 88.10 $\pm$ 0.13
TGAT-trunc 96.75 $\pm$ 0.02 95.77 $\pm$ 0.21 94.64 $\pm$ 0.01 98.78 $\pm$ 0.01 81.53 $\pm$ 0.14 84.10 $\pm$ 0.05
TGAT-unif 92.94 $\pm$ 0.15 96.33 $\pm$ 0.09 93.32 $\pm$ 0.02 98.75 $\pm$ 0.01 82.17 $\pm$ 0.08 82.21 $\pm$ 0.05
APAN-trunc 95.18 $\pm$ 0.91 96.02 $\pm$ 2.06 97.99 $\pm$ 0.03 CPU OOM 68.60 $\pm$ 2.41 80.47 $\pm$ 0.35
APAN-unif 95.73 $\pm$ 0.77 96.03 $\pm$ 1.31 96.55 $\pm$ 0.31 CPU OOM 78.64 $\pm$ 2.66 88.82 $\pm$ 0.28
@@
Task Method Wikipedia Reddit GDELT Ubuntu Wiki-talk
Transductive TGN-trunc 60.28 $\pm$ 4.09 74.38 $\pm$ 1.72 75.78 $\pm$ 8.67 20.94 $\pm$ 6.72 31.00 $\pm$ 5.02
TGN-unif 52.05 $\pm$ 4.65 65.36 $\pm$ 4.49 71.78 $\pm$ 0.16 21.87 $\pm$ 3.60 22.72 $\pm$ 6.55
TGAT-trunc 45.91 $\pm$ 1.15 60.19 $\pm$ 1.05 79.11 $\pm$ 0.04 22.37 $\pm$ 0.97 30.01 $\pm$ 0.84
TGAT-unif 23.17 $\pm$ 0.46 45.90 $\pm$ 0.47 74.89 $\pm$ 0.05 16.61 $\pm$ 0.44 18.01 $\pm$ 0.43
APAN-trunc 26.51 $\pm$ 1.46 43.68 $\pm$ 5.49 74.49 $\pm$ 4.14 5.79 $\pm$ 1.81 18.91 $\pm$ 9.17
APAN-unif 30.59 $\pm$ 3.51 36.29 $\pm$ 10.12 72.53 $\pm$ 0.22 11.60 $\pm$ 4.22 9.99 $\pm$ 1.77
NAT 77.51 $\pm$ 22.83 90.40 $\pm$ 4.63 GPU OOM 46.38 $\pm$ 9.14 74.17 $\pm$ 28.67
NAT-node 39.58 $\pm$ 6.21 39.89 $\pm$ 8.27 GPU OOM 18.93 $\pm$ 3.10 21.32 $\pm$ 3.59
NLB-edge 61.90 $\pm$ 2.31 75.40 $\pm$ 1.85 84.45 $\pm$ 0.84 24.34 $\pm$ 1.89 32.57 $\pm$ 1.49
NLB-node 54.42 $\pm$ 3.08 67.65 $\pm$ 5.31 76.60 $\pm$ 8.43 25.43 $\pm$ 0.97 31.95 $\pm$ 2.64
Inductive TGN-trunc 54.29 $\pm$ 3.13 58.20 $\pm$ 7.35 74.37 $\pm$ 1.26 19.58 $\pm$ 2.99 27.93 $\pm$ 2.94
TGN-unif 44.26 $\pm$ 2.32 53.02 $\pm$ 2.04 81.69 $\pm$ 3.73 18.14 $\pm$ 9.87 25.84 $\pm$ 0.68
TGAT-trunc 45.18 $\pm$ 1.56 50.67 $\pm$ 5.39 59.90 $\pm$ 0.59 22.41 $\pm$ 0.58 26.15 $\pm$ 1.50
TGAT-unif 15.60 $\pm$ 1.39 21.14 $\pm$ 6.09 51.72 $\pm$ 1.06 12.02 $\pm$ 1.03 20.51 $\pm$ 1.01
APAN-trunc 18.92 $\pm$ 0.28 34.85 $\pm$ 4.38 77.60 $\pm$ 0.23 7.16 $\pm$ 1.01 15.56 $\pm$ 3.52
APAN-unif 19.83 $\pm$ 1.77 38.15 $\pm$ 4.45 49.87 $\pm$ 3.95 11.63 $\pm$ 9.75 17.46 $\pm$ 5.45
NAT 75.02 $\pm$ 22.01 89.53 $\pm$ 4.08 GPU OOM 37.81 $\pm$ 12.40 39.02 $\pm$ 24.29
NAT-node 34.56 $\pm$ 4.18 50.58 $\pm$ 9.29 GPU OOM 19.64 $\pm$ 4.43 19.60 $\pm$ 6.73
NLB-edge 49.94 $\pm$ 2.12 61.74 $\pm$ 1.85 81.60 $\pm$ 4.17 21.13 $\pm$ 2.44 24.26 $\pm$ 4.60
NLB-node 44.59 $\pm$ 6.36 53.43 $\pm$ 8.03 71.58 $\pm$ 14.65 21.46 $\pm$ 2.97 25.39 $\pm$ 1.87
[…]
@@
  • Wikipedia [^2] http://snap.stanford.edu/jodie/wikipedia.csv [50, 19] records edit events on wiki pages. It is a bipartite graph where a set of nodes represents the editors and another set represents the wiki pages. The stamped link represents the edit events. The edge features are extracted from the contents of wiki pages. The node labels are binary which indicates whether a user is banned from posting. The original data dump is under the Creative Commons Attribution-Share-Alike 3.0 License and the Terms of Use is on the website (https://dumps.wikimedia.org/legal.html). The derived dataset by Kumar et al. [19] is open-sourced here.
@@
  • Reddit333 http://snap.stanford.edu/jodie/reddit.csv [51, 19] is a dataset of the post events by users on subreddits. It is also an attributed bipartite graph between users and subreddits. The node labels are also binary values that indicate whether a user is banned from posting to a subreddit. The original dataset has the Terms of Use in the website (https://pushshift.io/signup). The derived dataset by Kumar et al. [19] is open-sourced here.